OpenStudy (anonymous):

Find the surface area of a cone with a slant height of 19 and a radius of 7. Use 3.14 for π. Answer 569 570 571 572

OpenStudy (anonymous):

Surface Area of Cone = \[\left(\pi \times r \times l \right) + \left( \pi \times r ^{2} \right)\] Where r = radius and l = slant height. So the surface area = (3.14 x 7 x 19) + (3.14 x 49) = 571.48 = 571 (to the nearest whole number)